Let S be the set of real values of parameter k for which te eqn f(x) = 2x 3 - 3(2+k)x 2 +12kx has exactly one local minimum and exactly one local maximum.  S is a subset of

(a) (-4, ∞)      (b) (-3,3)        (c) (3,∞)   (d)  (- ∞,0

Grade:12

1 Answers

Ramesh V
14 years ago

f '(x) = 6(x2 - x(2+k) +2k ) =0

so roots of f '(x) are 2,k

f "(x) = 6( 2x -(2+k))

to have one local minima ( f "(x) > 0 )and local maxima ( f "(x) < 0 )

so , its k > 2 which is subset of (-4, infinity)

option A

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch